OpenStudy (anonymous):

3 − 5(x + 2) = 3(x + 1) + 6

OpenStudy (anonymous):

3-10x=3x+6 3-3=0 6-3=3 10x=3x+3 10x-3x=7x 3x-3x=0 7x=3 7x/3=2x 3/3=0 answer=2x or if you divide 3/7 it'll be just x